A bone is found in your back yard. You decide to d
Answer #1

Let z be the age of the bone

1Bq=60 disintegrations per minute

Then 0.097Bq=5.82 disintegrations per minute

And 0.23Bq=13.8 disintegrations per minute

(5.82/13.8)=(1/2)(z/5730 )       (the half life for carbon -14 is 5.73x103)

log(5.82/13.8)=(z/5730)log(1/2)

(log(0.42)/log(0.5))=(z/5730)

1.25x5730=z

z=7171 years